OpenStudy (anonymous):

The books in a private library are classified as fiction and nonfiction. There are 400 books in the library. There are 40 more fiction books than nonfiction books. Audrey randomly picks a book. A few minutes later, Ryan randomly picks one of the remaining books. What is the probability that both pick nonfiction books?

OpenStudy (anonymous):

180 × 189 400 × 400. 180 × 179 400 × 399. 180 × 179 400 × 400. 180 × 189 400 × 399.

OpenStudy (anonymous):

help me please

OpenStudy (texaschic101):

f + n = 400 n = f - 40 f + f - 40 = 400 2f = 440 f = 220 n = f - 40 n = 220 - 40 n = 180 f + n = 400 220 + 180 = 400 400 = 400 (correct) so there is a total of 220 fiction and 180 non-fiction 180/400 * 179/400

OpenStudy (anonymous):

lol thank you whats the answer

OpenStudy (anonymous):

ion think 180/400*179/400 is one of my choices

OpenStudy (anonymous):

haha thank you texas chick it is one of my choices i over looked it how do yu give metals